OpenStudy (anonymous):

Solve the Logarithmic equation. Please show me step by step, thank you! (Equation below)

OpenStudy (anonymous):

\[\log_{10}(x-30) = 3 - \log_{10}x \]

OpenStudy (anonymous):

@Mertsj perhaps you could be of assistance?

OpenStudy (mathstudent55):

Since the logs are of base 10, I'll just write log(x) log(x - 30) = 3 - log(x)

OpenStudy (mathstudent55):

Add log(x) to both sides

OpenStudy (mathstudent55):

log(x - 30) + log(x) = 3

OpenStudy (mathstudent55):

Do you know the law of the sum of logs?

OpenStudy (anonymous):

That's where if they're added you change it to multiplication, and if it's subtraction you change it to division, correct?

OpenStudy (mathstudent55):

Correct. What can you do to the left side now?

OpenStudy (anonymous):

log(x - 30)/x ?

OpenStudy (mathstudent55):

It's a sum of logs on the left, not a subtraction.

OpenStudy (anonymous):

Oh shoot, right, my bad, so then: log(x)(x - 30)

OpenStudy (mathstudent55):

log(x^2 - 30x) = 3

OpenStudy (anonymous):

Then I convert it to exponential? (x^2 - 30x) = 10^3

OpenStudy (mathstudent55):

Correct.

OpenStudy (mathstudent55):

Now solve the quadratic equation for x.

OpenStudy (anonymous):

Thank you very much!

OpenStudy (mathstudent55):

wlcm

OpenStudy (mathstudent55):

remember to check both solutions of the quadratic eq in the original logarithmic eq since the domain of the log function is restricted.

OpenStudy (anonymous):

Cannot be negative log or log(0) correct?

OpenStudy (mathstudent55):

Correct. log(x) is defined for positive x only. That does not mean that a negative value of x from the solution of the quadratic eq will produce an invalid argument for the log function. The best way to find out is to check both solutions. In other words, a negative solution to the quadratic eq may still work. You need to check both solutions to be sure.

OpenStudy (anonymous):

Okay. and since I've got your attention still, what would I do for the next step in this problem? It's the same kind, except it was subtraction, so it was changed to division. Still not good at these types of problems. :/ \[\log_{2} (x+2)/(x-5) = 1\]

OpenStudy (mathstudent55):

You mean the log base 2 of the entire fraction? Or the log base 2 of (x + 2), then divided by x - 5?

OpenStudy (anonymous):

Yes, log base 2 then in parenthesis the x+2 over x-5

OpenStudy (mathstudent55):

OK, then change the log eq into an exponential eq

OpenStudy (anonymous):

Ah, alright, got it. Thanks again!